The sum of three numbers is -2. The sum of three times the first number, twice the second number, and the third number is 9. The difference between the second number and half the third number is 10. Find the numbers.
Answer by CubeyThePenguin(3113)

x + y + z = -2
3x + 2y + z = 9
y - z/2 = 10 ----> 2y - z = 20

Eliminate the variable z.

equation 2 and equation 3:
3x + 2y + z = 9
2y - z = 20
-----------------------
3x + 4y = 29

equation 1 and equation 3:
x + y + z = -2
2y - z = 20
---------------------
x + 3y = 18

Use any method (elimination, substitution, etc) to solve for x and y. Then, substitute those values to solve for z.